OpenStudy (anonymous):

find the slop of the line x-5y=8.

OpenStudy (anonymous):

Solve for y, and the coefficient of x will be the slope. This is because \(y = mx + b\).

OpenStudy (anonymous):

-5y=8-x y=(8-x)/(-5) y=-8/5+x/5 y=(1/5)x-8/5 since y=mx+b, and m is the slope ...

OpenStudy (anonymous):

so the answer is 1/5 correct ?

OpenStudy (texaschic101):

x - 5y = 8 -5y = -x + 8 y = 1/5x - 8/5 yes...the slope is 1/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you Leon88 and texaschic101

OpenStudy (texaschic101):

anytime :) Leon should get the medal...he answered first

OpenStudy (anonymous):

indeed I did. Thank you

OpenStudy (anonymous):

Actually, I answered first.

OpenStudy (texaschic101):

thats true...my bad...give it to bloopman

OpenStudy (texaschic101):

there you go bloopman...you got one also :)

OpenStudy (anonymous):

Ah, no worries. Thanks anyhow.

OpenStudy (texaschic101):

you deserve it :)